If 23.9 mL of the strontium hydroxide solution was
Answer #1

Sr(OH)2 (aq.)+2 HNO3 (aq.) ---------> Sr(NO3)2 (aq.) + 2 H2O(l)

1 mol Sr(OH)2 requires 2 mol of HNO3

Fromula,

M1V1 / n1 = M2V2 / n2

M1 * 31.5 / 2 = M2 * 23.9 / 1

M1 = 1.52 * M2

COncentration of acid i.e HNO3 must be 1.52 times to concetration of base i.e Sr(OH)2.